The parallel sides of a trapezoid are 8 inches and 12 inches long, while one of the non- parallel sides is 6 inches long. How far must this side be extended to meet the extension of the opposite side? What are the possible lengths for the opposite side?
Answer by mananth(16946)

.
AB ||CD
Therefore triangle EAB ~ triangle ECD ( Two angles congruent)
Therefore AB/CD = EA/EC
8/12 =x/(x+6)
3/4 = x/(x+6)
3x+18=4x
x=18